# MoIRA: Modular Instruction Routing Architecture for Multi-Task Robotics

### v1

@@ Abstract
Mixture-of-Experts (MoE) approaches have recently gained traction in robotics applications due to their ability to dynamically allocate computational resources and specialize sub-networks for distinct tasks or environmental contexts, enabling more efficient decision-making. Such systems often comprise sparsely activated experts combined under a single monolithic architecture and require a well-configured internal routing mechanism, which does not allow for selective low-level expert and router customization and requires additional training. We propose MoIRA, an architecture-agnostic modular MoE framework designed to coordinate existing experts with an external text-based router. MoIRA incorporates two zero-shot routing options: embedding-based similarity and prompt-driven language model inference. In our experiments, we choose large Vision-Language-Action models, gr00t-N1 and $\pi_{0}$, as the underlying experts, and train low-rank adapters for low-overhead inference. We evaluate MoIRA on various GR1 Humanoid tasks and LIBERO Spatial and Goal benchmarks, where it consistently outperforms generalist models and competes with other MoE pipelines. Additionally, we analyse the robustness of the proposed approach to the variations of the instructions. While relying solely on textual descriptions of tasks and experts, MoIRA demonstrates the practical viability of modular deployment with precise, low-effort routing and provides an alternative, scalable foundation for future multi-expert robotic systems.
Keywords: Vision-Language-Action models; Modular Policies; Robotic Control; Mixture-of-Experts; Multi-Task Learning; Expert Routing.
@@ 1 Introduction
Figure 1: Overview of our MoIRA framework: (1) Fine-tuned agents, each a VLA with a dedicated LoRA adapter; (2) Meta‐descriptions, concise summaries $m_{i}$ characterising each agent’s expertise; and (3) Router core, which infers the specialist index $r$ via one of our two proposed strategies: embedding-based similarity matching (via text encoder $E$) between task $t$ and meta-descriptions $m_{i}$, or prompt-driven language model routing. The resulting index $r$ directs the selected specialist for downstream policy inference.
Concurrently, Mixture-of-Experts (MoE) architectures have regained interest as a modular approach to specialization. Originally proposed by Jacobs et al. [14] for adaptive task decomposition, MoEs are now widely explored for improving efficiency in large language models [6] and in robotic learning [12, 24, 29, 28, 13].
In dexterous manipulation, residual MoE structures [13] enable the composition of expert policies without relying on language inputs or transformers, using fixed routing strategies across object-centric subtasks. MoLe-VLA [28] introduces a spatial-temporal aware router (STAR) to dynamically activate sublayers in VLA models based on token relevance and task structure.
In the locomotion domain, MoRE [29] applies sparse expert activation to quadruped control, routing tasks through a shared vision-language interface to balance sample efficiency with generalization. Similarly, GERM [24] demonstrates expert-based generalist control via MoEs for legged agents.
Additional work explores MoE strategies for policy inference efficiency. MoDE [22] leverages sparse expert denoisers in diffusion transformer policies for LIBERO tasks, while Tra-MoE [27] builds on Any-Point Trajectory Modeling [26] to improve multi-domain policy prediction via sparsely-gated transformers.
@@
While these systems show promise, they share key limitations. Generalist models trained on broad data tend to suffer from degraded per-task performance and high runtime memory usage. Expert-based MoEs require orchestrated training and internal routing mechanisms, often tied to monolithic model structures. This introduces a tradeoff between specialization, modularity, and deployment flexibility that has not been fully addressed.
@@
To address these limitations, we adopt a modular architectural perspective wherein each expert can be independently developed, customized, and optimized. This decouples training from deployment, and allows for flexible reuse across tasks, embodiments, and input modalities. We therefore propose the Modular Instruction Routing Architecture, MoIRA (Figure 1), a framework designed to perform zero-shot model routing via natural language task and expert descriptions. MoIRA sidesteps the scalability constraints of monolithic MoEs by leveraging a pool of pretrained specialists, each finetuned on a focused domain. A lightweight meta-controller dynamically selects the most relevant expert, using either embedding-based similarity or prompt-driven reasoning over the textual task specification.
@@
We evaluate MoIRA on two robotic benchmarks: GR1 [20], which covers embodiment variation (full-body, arms-only, and arms-with-waist), and LIBERO [18], which separates tasks into Goal and Spatial semantic categories. For these experiments, we instantiate MoIRA with the GR00t-N1 and $\pi_{0}$ VLA backbones, using LoRA adapters [11] to enable efficient specialist training. The routing module is pretrained and frozen, requiring no additional tuning to map tasks to experts.
@@
    We evaluate two routing strategies – cosine similarity with MiniLM [21]) and prompt-based inference with SmolLM2-1.7B [1] – and demonstrate robustness under perturbed inputs.
@@
    We validate MoIRA on the GR1 and LIBERO benchmarks, showing that it consistently outperforms generalist models and other MoE approaches on target tasks and previously unseen tasks.
@@
    We provide an empirical analysis of runtime memory usage and LoRA adapter switching latency, demonstrating the feasibility of multi-expert deployment under hardware constraints.
By decoupling task semantics from execution, MoIRA enables scalable, modular control. It offers a practical step toward robotic agents that generalize across tasks while benefiting from specialization without relying on monolithic training pipelines.
@@ 2 Related Work
Classical approaches in robotics rely on RL and IL, both of which have shown strong performance in simulation and real-world environments. RL methods perform well under dense, shaped reward functions, using techniques such as off-policy replay or on-policy actor-critic optimization [23, 7]. However, they often fail in sparse reward regimes due to delayed feedback and exploration bottlenecks. IL methods, particularly behavior cloning [17], bypass reward engineering by mimicking expert demonstrations, yet suffer from distributional shift and limited generalization when faced with unseen states or perturbations.
Foundation models for VLA control. The emergence of foundation vision-language models (VLMs) has shifted the paradigm in robotic perception and planning. Models such as PaLI-Gemma [25], LLaVA [19], and Qwen-VL [2] demonstrate impressive multimodal grounding capabilities, allowing them to parse natural language prompts and interpret visual scenes. While not trained for control, these models have inspired a new class of instruction-conditioned control architectures.
@@
To bridge the gap between perception and action, recent work extends VLMs into vision-language-action (VLA) models. These architectures integrate visual-language encoders with control heads, supporting end-to-end policy learning from natural language instructions. Examples include RT-2 [5] and the RT-X family [8], which are trained on large-scale robot datasets and achieve generalization across tasks. OpenVLA [15] and MiniVLA [3] focus on smaller, instruction-following backbones, while $\pi_{0}$ [4] and GR00t-N1 [20] support scalable deployment through modular and embodiment-aware control heads. These models commonly use datasets like Open-X Embodiment (OXE) [8] and Ego4D [9] to align multimodal inputs with action trajectories.
While VLAs are powerful generalists, their monolithic design often results in inefficiencies when scaled to diverse task domains. They may suffer from overfitting to frequent patterns, struggle with rare or fine-grained instructions, and require extensive resources during inference. These limitations motivate the development of modular alternatives that preserve generalization while enabling efficient task-level adaptation.
Figure 2: Example rollouts from tabletop manipulation tasks using the GR1 Humanoid embodiment and the LIBERO Spatial benchmark with a Panda Franka arm.
Modularity and mixture-of-experts in robotic control. Mixture-of-Experts (MoE) frameworks provide a natural structure for modularization. Initially proposed for adaptive function decomposition [14], MoEs have been extensively adopted in large-scale NLP and are now gaining traction in robotics for task specialization and efficiency [6]. Their appeal lies in sparse expert activation, which allows scaling while keeping inference efficient.
In robotic manipulation, residual MoE architectures [13] demonstrate how multiple base policies can be composed without requiring language conditioning or transformers. These models use fixed routing strategies over handcrafted task splits and achieve strong results in structured environments. MoLe-VLA [28] introduces a learned Spatial-Temporal Aware Router (STAR) that dynamically activates sublayers within a VLA backbone, reducing unnecessary computation while preserving output quality.
For locomotion, Zhao et al. propose MoRE [29], which integrates sparse expert activation into a quadruped control policy using a shared VLM-based interface. Similarly, GERM [24] leverages MoE modules to build scalable, generalist locomotion agents that maintain specialist precision. These works highlight how modularity can be applied across different embodiment types, even in low-level control settings.
Recent work has also explored MoE routing for efficient policy inference. MoDE [22] introduces a diffusion-based transformer model with sparsely activated expert denoisers, targeting sample-efficient multitask learning in manipulation. Tra-MoE [27] builds on Any-point Trajectory Modeling [26], employing a sparsely-gated transformer to improve cross-domain generalization in imitation learning. Both designs rely on internal routing mechanisms and integrated training pipelines, which limit post-hoc adaptability and expert reuse.
Figure 3: Serving and inference configurations in MoIRA. (I) All-agents-in-memory: Each specialist, identified by a unique LoRA adapter ID, is kept resident in GPU memory. This setup enables near-instant agent switching but requires significantly more VRAM, scaling with the number of specialists. (II) Dynamic LoRA adapter loading: A single shared backbone dynamically loads the required LoRA adapter based on router input (e.g., ”Full upper body specialist (ID: 2)”). This configuration reduces memory overhead to one model checkpoint and a single adapter, but introduces up to 9s latency when swapping specialists.
Positioning MoIRA. These prior systems demonstrate the benefits of modularity but also expose critical tradeoffs. Most MoE-based control systems require tightly coupled routing logic, joint training, or customized architectures that reduce interpretability and limit plug-and-play integration. In contrast, MoIRA adopts an external routing design, treating expert selection as an independent classification task driven by pretrained language models. Its use of LoRA adapters enables low-overhead specialization, and its modular layout supports architectural flexibility across tasks and embodiments. By decoupling routing from execution, MoIRA enables expert policies to be updated or replaced independently. Unlike internal token-level gating, its episodic instruction-based routing supports fast switching, simplified deployment, and interpretability of decision boundaries. These design choices make MoIRA especially suited for real-world deployments where modularity, memory efficiency, and runtime specialization are essential.
@@ 3 Methods
At inference time, MoIRA performs two steps: routing and expert execution (Algorithm 1). Given a task instruction $t$ and expert descriptions $\{m_{i}\}_{i=1}^{K}$, it selects the expert index $r$ using either cosine similarity over pretrained text embeddings:
@@
or prompt-based reasoning via a frozen LM with Chain-of-Thought formatting:
@@
The selected expert $\mathcal{E}_{r}$ is then executed either directly from memory or by loading LoRA adapter weights to produce the output trajectory $\tau$ and result $o$.
Algorithm 1 MoIRA Routing and Execution
@@ 3.1 Benchmarks and Datasets
    GR1: We use the GR1 benchmark containing three embodiments (arms-only, arms & waist, and full upper body) of GR1 humanoid. Each task contains an annotated task instruction retrieved from tasks.jsonl metadata file.
@@
    LIBERO: LIBERO benchmark tasks are extracted from .bddl files. We use a modified version of LIBERO splits from [16] and select Spatial and Goal task categories of the benchmark as our semantic axis.
@@ 3.2 Specialist Finetuning
Figure 5: The initial stage of MoIRA’s agentic adapter fine-tuning. We finetune Vision-Language-Action backbones (Gr00t-N1, $\pi_{0}$-base) and derive modular specialists via LoRA: embodiment adapters for GR1 manipulation tasks and semantic adapters for LIBERO goal/spatial tasks. Although demonstrated on VLAs, MoIRA’s unified interface can accommodate any agentic backbone, e.g. transformer-based or model-based approaches.
For GR1, a dedicated GR00t-N1-2B specialist is LoRA fine-tuned for 5K steps for each embodiment. We select three representative tasks to capture variation in embodiment types, task dynamics, and complexity – Pouring, CanSort, and PlacematToPlate. In addition, we train a jointly-tuned generalist model across all three tasks and reserve WineToCabinet (arms & waist), CuttingboardToTieredBasket (arms & waist), and Coffee (full upper body) as previously unseen (hold-out) tasks for evaluation.
For LIBERO, we use the $\pi_{0}$-base-3.3B foundation VLA as the backbone. We LoRA-fine-tune three models: a spatial task expert, a goal task expert, and a generalist, jointly-trained on both tasks – each model tuned for 30K steps using default hyperparameters. The fine-tuning stage is described in Figure 5.

### v2

@@ Abstract
Mixture-of-Experts (MoE) approaches have gained traction in robotics for their ability to dynamically allocate resources and specialize sub-networks. However, such systems typically rely on monolithic architectures with rigid, learned internal routing, which prevents selective expert customization and necessitates expensive joint training. We propose MoIRA, an architecture-agnostic modular framework that coordinates decoupled experts via an external, zero-shot text router. MoIRA employs two routing strategies: embedding-based similarity and prompt-driven language model inference. Leveraging Gr00t-N1 and $\pi_{0}$ Vision-Language-Action models with low-rank adapters, we evaluate MoIRA on GR1 Humanoid tasks and LIBERO benchmarks. Our approach consistently outperforms generalist models and competes with fully trained MoE pipelines. Furthermore, we demonstrate system robustness against instruction perturbations. By relying on textual descriptions for zero-shot orchestration, MoIRA proves the viability of modular deployment and offers a scalable, flexible foundation for multi-expert robotic systems.
keywords:
Vision-Language-Action models, Modular Policies, Robotic Control, Mixture-of-Experts, Multi-Task Learning, Expert Routing
journal: Neurocomputing
[…]
[label1]organization=Department of Multimedia Systems, National University of Kyiv-Mohyla Academy,addressline=Hryhoriya Skovorody St, 2, city=Kyiv, postcode=04655, country=Ukraine
[…]
[label2]organization=Department of Mathematics, National University of Kyiv-Mohyla Academy,addressline=Hryhoriya Skovorody St, 2, city=Kyiv, postcode=04655, country=Ukraine
footnotetext: This manuscript is the author-accepted version of an article published in Neurocomputing, vol. 674, 132962 (2026). The final version of record is available at https://doi.org/10.1016/j.neucom.2026.132962.
@@ 1 Introduction
Figure 1: Overview of the MoIRA framework. The architecture decouples policy learning from task assignment by coordinating (1) a pool of independently fine-tuned specialist agents via (2) textual meta-descriptions. The (3) Router core performs zero-shot assignment of the unseen task $t$ to the optimal specialist index $r$ using either embedding similarity or prompt-driven reasoning. This design allows modular expert addition without retraining the routing mechanism.
Concurrently, Mixture-of-Experts (MoE) architectures have regained interest as a modular approach to specialization. Originally proposed by Jacobs et al. [13] for adaptive task decomposition, MoEs are now widely explored for improving efficiency in large language models [6] and in robotic learning. In dexterous manipulation, residual MoE structures [12] enable the composition of expert policies without relying on language inputs or transformers. MoLe-VLA [29] introduces a spatial-temporal aware router to dynamically activate sublayers in VLA models based on token relevance. In the locomotion domain, MoRE [30] and GERM [24] apply sparse expert activation to quadruped control, balancing sample efficiency with generalization. Similarly, recent inference-optimization works like MoDE [21] and Tra-MoE [28] leverage sparsely-gated transformers to improve multi-domain policy prediction.
@@
While these systems show promise, they share key limitations related to their architectural rigidity. Generalist models trained on broad data tend to suffer from degraded per-task performance and high runtime memory usage. Conversely, existing expert-based MoEs typically require orchestrated training and internal routing mechanisms tied to specific monolithic model structures. This introduces a critical tradeoff between specialization, modularity, and deployment flexibility that has not been fully addressed.
@@
To address these limitations, we adopt a modular architectural perspective wherein each expert can be independently developed, customized, and optimized. This decouples training from deployment and allows for flexible reuse across tasks and embodiments. We therefore propose Modular Instruction Routing Architecture, MoIRA (Figure 1), a framework designed to perform zero-shot episodic model routing via natural language task and expert descriptions. Beyond routing, MoIRA is designed for practical specialist serving via adapter-based experts, supporting both disk-based swapping and multi-adapter [19, 23] hot-switching for low-latency deployment. MoIRA sidesteps the scalability constraints of monolithic MoEs by leveraging a pool of pretrained specialists, each finetuned on a focused domain. A lightweight meta-controller dynamically selects the most relevant expert using either embedding-based similarity or prompt-driven reasoning over the textual task specification.
@@
We evaluate MoIRA on two robotic benchmarks: GR1 [18], which covers embodiment variation (full-body, arms-only, and arms & waist), and LIBERO [16], which separates tasks into Goal and Spatial semantic categories. For these experiments, we instantiate MoIRA with the GR00t-N1 and $\pi_{0}$ VLA backbones, using LoRA adapters [11] to enable efficient specialist training. The routing module is pretrained and frozen, requiring no additional tuning to map tasks to experts.
@@
    We evaluate two routing strategies – cosine similarity with MiniLM [20] and prompt-based inference with SmolLM2-1.7B [1] – and demonstrate robustness under perturbed inputs.
@@
    We validate MoIRA on the GR1 and LIBERO benchmarks, showing that it consistently outperforms or achieves parity with generalist models and other MoE approaches on target tasks and previously unseen tasks.
@@
    We provide an empirical analysis of inference-time expert serving, quantifying the trade-off between VRAM usage and switching latency across (i) fully instantiated adapters, (ii) disk-based swapping, and (iii) multi-LoRA hot-switching for scalable multi-expert deployment.
By decoupling task semantics from execution, MoIRA enables scalable, modular control. It utilizes an architecture-agnostic, external routing mechanism to coordinate a dynamic pool of specialized experts, each implemented as a lightweight LoRA adapter. The zero-shot, language-based routing allows for the independent addition, update, or replacement of experts without costly joint training or retraining the router. It validates a flexible design paradigm for robotic agents that generalize across tasks while benefiting from specialization, providing an alternative to monolithic training pipelines.
Figure 2: Visualization of policy rollouts. GR1 Humanoid embodiment (arms-only, full upper body) completing pouring and pick-and-place tasks and a Panda Franka arm successfully executing manipulation primitives on the LIBERO Spatial benchmark.
@@ 2 Related Work
Policy learning for robotic manipulation. Classical approaches in robotics rely on RL and IL, both of which have shown strong performance in simulation and real-world environments. RL methods perform well under dense, shaped reward functions, using techniques such as off-policy replay or on-policy actor-critic optimization [22, 7]. However, they often fail in sparse reward regimes due to delayed feedback and exploration bottlenecks. IL methods, particularly behavior cloning [15], bypass reward engineering by mimicking expert demonstrations, yet suffer from distributional shift and limited generalization when faced with unseen states or perturbations.
@@
Foundation models for VLA control. The emergence of VLMs has shifted the paradigm in robotic perception and planning. Models such as PaLI-Gemma [25], LLaVA [17], and Qwen-VL [2] demonstrate impressive multimodal grounding capabilities. While not trained for control, these models have inspired a new class of instruction-conditioned control architectures. To bridge the gap between perception and action, recent work extends VLMs into VLA models. These architectures integrate visual-language encoders with control heads, supporting end-to-end policy learning from natural language instructions. Examples include RT-2 [5] and the RT-X family [8]. OpenVLA [14] and MiniVLA [3] focus on smaller, instruction-following backbones, while Physical Intelligence $\pi_{0}$ [4] and NVIDIA Gr00t-N1 [18] support scalable deployment through modular control heads. While VLAs are powerful generalists, their monolithic design often results in inefficiencies when scaled to diverse task domains, motivating the development of modular alternatives that preserve generalization while enabling efficient task-level adaptation.
Figure 3: Comparison of expert serving strategies in MoIRA. We illustrate three deployment methods: (I) Fully Instantiated maintains active copies of all experts in memory; while fast, this incurs prohibitive linear VRAM scaling. (II) Disk-Based Swapping reduces memory footprint by loading adapters from storage on-demand, but introduces a high I/O latency bottleneck ($\sim$10s). (III) Multi-LoRA Serving keeps the backbone resident and maintains multiple lightweight adapters concurrently, enabling hot-switching via pointer changes ($\sim$20ms) while avoiding full model replication and large I/O swap costs.
Modularity and Mixture-of-Experts in Robotics. MoE frameworks provide a natural structure for modularization [13, 6]. However, current robotic MoEs typically rely on internal routing mechanisms that are tightly coupled to the model architecture. For instance, MoLe-VLA [29] utilizes a learned Spatial-Temporal Aware Router for dynamic layer skipping, while MoRE [30] and GERM [24] employ sparse expert activation within shared transformer backbones for locomotion. Unlike these systems, which require joint training and specific internal gating structures, MoIRA treats experts as independent black-box modules, enabling architecture-agnostic coordination.
Efficient Inference and Parameter Merging. Recent work has focused on reducing the computational overhead of MoEs. MoDE [21] and Tra-MoE [28] leverage sparsely-gated transformers to optimize multitask policy prediction. Beyond sparse activation, He et al. [10] demonstrate that merging expert parameters prior to inference can significantly reduce FLOPs and latency without maintaining discrete active experts. While parameter merging effectively lowers computational cost, it assumes that experts share a merge-compatible architecture and does not address the VRAM constraints of hosting diverse, non-mergeable specialist models. MoIRA addresses this distinct hardware constraint via dynamic adapter swapping (Figure 3), trading IO latency for the ability to deploy a vast, modular library of LoRA specialists, constrained by disk space rather than VRAM, supporting episodic multi-task deployment on high-capacity hardware.
Complementary to parameter merging, recent systems work focuses on serving large numbers of LoRA adapters efficiently by keeping a shared backbone resident while multiplexing many low-rank adapters with optimized kernels. Frameworks such as S-LoRA [23] and LoRAX [19] enable multi-adapter concurrency and millisecond-scale activation, directly addressing the deployment bottleneck in modular expert libraries. MoIRA leverages this multi-LoRA operating point to expand beyond episodic disk swapping while preserving explicit expert boundaries.
Semantic Routing and Positioning MoIRA. Our approach to expert orchestration shares strong methodological parallels with ”divide-and-conquer” strategies in NLP, most notably the work of Wang et al. [26]. While their framework was developed to decouple data into semantic clusters for zero-shot dialogue state tracking, MoIRA establishes an analogous architecture for the domain of robotic control. Wang et al. apply this separation of concerns to abstract dialogue slots, while MoIRA validates the paradigm’s efficacy in handling heterogeneous embodiment and physical task variations. This cross-domain convergence highlights the fundamental robustness of decoupling routing from execution: it ensures that expert policies can be updated, replaced, or added independently without retraining the router, offering a scalable alternative to internal gating that is particularly proven for dynamic real-world environments.
@@ 3 Methods
At inference time, MoIRA performs two steps: routing and expert serving (Algorithm 1). Given a task instruction $t$ and expert descriptions $\{m_{i}\}_{i=1}^{K}$, it selects the expert index $r$ using either cosine similarity over pretrained text embeddings:
@@
or prompt-based reasoning via a frozen LM with Chain-of-Thought [27] formatting:
@@
The selected expert $\mathcal{E}_{r}$ is then served under one of several serving regimes (Figure 3) to produce the output trajectory $\tau$ and result $o$.
Algorithm 1 MoIRA Routing and Serving
@@
The computational complexity of the routing step scales linearly $O(K)$ with the number of experts $K$. Embedding router requires a single forward pass of the encoder $E(t)$ and $K$ dot-product operations, resulting in negligible latency. For the language model router, context length grows linearly with the length of meta-descriptions $\sum_{i=1}^{K}|m_{i}|$, with an empirical average description length of $\approx 35$ tokens observed in our experiments. While more semantically robust, inference latency increases with the size of the expert pool.
The primary system bottleneck lies in the expert swap stage. We analyze this constraint through three distinct serving configurations. The simplest, in-memory switching (i), maintains all specialists in VRAM, enabling instantaneous transitions but imposing prohibitive memory requirements that scale linearly with the expert pool. Conversely, disk-based swapping (ii) optimizes for memory by storing inactive experts on disk, though this introduces significant I/O latency during retrieval. To address this trade-off, we consider multi-LoRA serving (iii) by utilizing frameworks such as S-LoRA [23] and LoRAX [19] as an effective intermediate approach. By leveraging optimized kernels, this method allows for the dynamic swapping of VLA adapters (e.g., 4B parameters) in up to 20ms, effectively balancing low memory footprint with near-real-time responsiveness.
@@ 3.1 Benchmarks and Datasets
    GR1: We use the GR1 benchmark [18] containing three embodiments (arms-only, arms & waist, and full upper body) of GR1 humanoid. Each task contains an annotated task instruction retrieved from tasks.jsonl metadata file.
@@
    LIBERO: LIBERO benchmark [16] tasks are extracted from .bddl files. We use a modified version of LIBERO splits from [14] and select Spatial and Goal task categories of the benchmark as our semantic axis.
@@ 3.2 Specialist Finetuning
Figure 4: The initial stage of MoIRA’s agentic adapter fine-tuning. We finetune VLA backbones (Gr00t-N1, $\pi_{0}$-base) and derive modular specialists via LoRA: embodiment adapters for GR1 manipulation tasks and semantic adapters for LIBERO Goal and Spatial tasks. Although demonstrated on VLAs, MoIRA’s unified interface can accommodate any agentic backbone, e.g. transformer-based or model-based approaches.
For GR1, a dedicated GR00t-N1-2B specialist is LoRA fine-tuned for 5K steps for each embodiment. We select three representative tasks to capture variation in embodiment types, task dynamics, and complexity – Pouring, CanSort, and PlacematToPlate. In addition, we train a jointly-tuned generalist model across all three tasks and reserve WineToCabinet (arms & waist), CuttingboardToTieredBasket (arms & waist), and Coffee (full upper body) as previously unseen (held-out) tasks for evaluation.
For LIBERO, we use the $\pi_{0}$-base-3.3B foundation VLA as the backbone. We create LoRA-adapters for three experts: a spatial task expert, a goal task expert, and a generalist, jointly trained on both tasks – each model tuned for 30K steps using default hyperparameters. The fine-tuning stage is described in Figure 4.